# Online Smoothed Demand Management

### v1

@@ Appendix F Deferred Proofs from § 5
To solve for $\alpha^{\star}$, we can use Grönwall’s inequality [Mitrinovic:91, Theorem 1, p. 356]. First, observe that $h(x)$ must satisfy the following:
@@
By Grönwall’s inequality [Mitrinovic:91, Theorem 1, p. 356], it follows that:
@@
Recall that we have a boundary condition that $h(p_{\min})=1$ – we can combine this with the above to obtain:
@@
$\alpha^{\star}\geq\frac{\omega}{\left[W\left(-\frac{\left((1+c+\varepsilon)p_{\max}-(1+\varepsilon)p_{\min}\right)e^{\frac{-\omega\nicefrac{{2\kappa}}{{T}}-p_{\min}c-(1+c+\varepsilon)p_{\max}}{(1+c+\varepsilon)p_{\max}+2\kappa}}}{(1+c+\varepsilon)p_{\max}+2\kappa}\right)+\frac{(1+c+\varepsilon)p_{\max}+p_{\min}c-\omega\nicefrac{{2\kappa}}{{T}}}{(1+c+\varepsilon)p_{\max}+2\kappa}\right]},$
@@
Let $h(x)$ denote a conversion function that maps $p_{\min},p_{\max}]\to[0,1]$ – this describes ALG’s purchasing decisions before the final batch of prices ($p_{\max}$) in the instance. Since ALG is deterministic and purchasing is unidirectional (irrevocable), we must have that $h(x-\varrho)\geq h(x)$, i.e., $h(x)$ is non-increasing in $p_{\min},p_{\max}]$. Intuitively, the entire demand should be purchased before the end of the sequence if the lowest price appears, i.e., $h(p_{\min})=1$.
@@
By Grönwall’s inequality [Mitrinovic:91, Theorem 1, p. 356], it follows that:
@@
Recall that we have a boundary condition that $h(p_{\min})=1$ – we can combine this with the above to obtain:
@@
$\alpha_{\texttt{IF}}\geq\left[W\left(-\frac{\left(p_{\max}-p_{\min}\right)e^{-1}}{p_{\max}+\frac{2\kappa}{T(1+\varepsilon)}}\right)+1\right]^{-1}.$ (25)
@@
By Grönwall’s inequality [Mitrinovic:91, Theorem 1, p. 356], it follows that:
@@
Recall that we have a boundary condition that $h(p_{\min})=1$ – we can combine this with the above to obtain:
@@
$\alpha_{\texttt{IB}}\geq\left[W\left(-\frac{\left((1-c)p_{\max}-p_{\min}\right)e^{-\frac{(1+\varepsilon)p_{\max}+2\delta+\frac{2\gamma}{T}}{(1+\varepsilon)p_{\max}+2\kappa}}}{(1+\varepsilon)p_{\max}+2\kappa}\right)+\frac{(1+\varepsilon)p_{\max}+2\delta+\frac{2\gamma}{T}}{(1+\varepsilon)p_{\max}+2\kappa}\right]^{-1}.$ (26)
@@
Let $h(x)$ denote a conversion function that maps $p_{\min},p_{\max}]\to[0,1]$ – this describes ALG’s purchasing decisions before the final batch of prices ($p_{\max}$) in the instance. Since ALG is deterministic and purchasing is unidirectional (irrevocable), we must have that $h(x-\varrho)\geq h(x)$, i.e., $h(x)$ is non-increasing in $p_{\min},p_{\max}]$. Intuitively, the entire demand should be purchased before the end of the sequence if the lowest price appears, i.e., $h(p_{\min})=1$.
@@
By Grönwall’s inequality [Mitrinovic:91, Theorem 1, p. 356], it follows that:
@@
Recall that we have a boundary condition that $h(p_{\min})=1$ – we can combine this with the above to obtain:
@@
Due to the tracking target, any procurement before the second-to-last batch incurs a tracking penalty of $\eta h(x)$, since the target is zero during this time. Furthermore, during the last two batches, ALG incurs a tracking penalty of $\eta+\eta(1-h(x))$, since the target is $\nicefrac{{1}}{{m}}$ during the first $m$ prices in the batch, and zero during the second $m$ prices in the batch – note that ALG must purchase (1h(x))(1-h(x)) at the last price to satisfy the deadline. Thus, the total tracking penalty incurred by ALG is given by $2\eta$.
@@
To solve for $\alpha^{\star}$, we can use Grönwall’s inequality [Mitrinovic:91, Theorem 1, p. 356]. First, observe that $h(x)$ must satisfy the following:
@@
Recall that we have a boundary condition that $h(p_{\min})=1$ – we can combine this with the above to obtain:
@@
$\alpha^{\star}\geq\frac{\omega}{\left[W\left(-\frac{\left((1+c+\varepsilon)p_{\max}-(1+\varepsilon)p_{\min}\right)e^{\frac{-\omega\nicefrac{{2\delta}}{{T}}-p_{\min}c-(1+c+\varepsilon)p_{\max}}{(1+c+\varepsilon)p_{\max}+2(\eta+\delta)}}}{(1+c+\varepsilon)p_{\max}+2(\eta+\delta)}\right)+\frac{(1+c+\varepsilon)p_{\max}+p_{\min}c-\omega\nicefrac{{2\delta}}{{T}}}{(1+c+\varepsilon)p_{\max}+2(\eta+\delta)}\right]},$
@@
We now proceed to prove the worst-case competitive ratio of PALD– for notational brevity, the following considers any arbitrary $\mathcal{I}\in\Omega$ (i.e., any arbitrary instance of OSDM). We recall the definition of active and inactive periods from Definition E.2, and the additional notations relating to driver indexing used in the proof of Theorem 4.1 in Appendix E.
@@
$\texttt{PALD}(\mathcal{I})$ $\leq\sum_{i=1}^{n}\sum_{j=1}^{\nu_{i}}\int_{0}^{\hat{w}_{b}^{(i,j)}}\kern-10.00002pt\hat{\phi}_{b}^{(i,j)}(u)du+\sum_{\tau=1}^{T}\left(\int_{0}^{\hat{w}_{f}^{(\tau)}}\kern-10.00002pt\hat{\phi}_{f}^{(\tau)}(u)du+\int_{0}^{\hat{v}_{f}^{(\tau)}}\kern-10.00002pt\hat{\psi}^{(\tau)}(z)dz\right)$ (27)
$\quad+\left(D_{b}+D_{f}-\sum_{i=1}^{n}\sum_{j=1}^{\nu_{i}}\hat{w}_{b}^{(i,j)}-\sum_{\tau=1}^{T}\hat{w}_{f}^{(\tau)}\right)(p_{\max}+2\gamma)$ (28)
$\quad+D_{b}\left(\bar{p}(c+\varepsilon)+2\delta\right)+\left(D_{f}-\sum_{\tau=1}^{T}\hat{v}_{f}^{(\tau)}\right)(p_{\max}(c+\varepsilon)+2\delta)$ (29)
$\quad-cp_{\min}\left(\sum_{i=1}^{n}\sum_{j=1}^{\nu_{i}}\hat{w}_{b}^{(i,j)}+\sum_{\tau=1}^{T}\hat{w}_{f}^{(\tau)}\right)+\hat{s}p_{\max}.$ (30)
@@
The proof follows by the same logic as Lemma E.4 – the main change is with respect to (27), to capture the fact that the threshold functions used by PALD are the learned variants, denoted by $\hat{\phi}_{b}$ and $(\hat{\phi}_{f},\hat{\psi}_{f})$, instead of the analytical variants $\phi_{b}$ and $(\phi_{f},\psi_{f})$ used by PAAD.
@@
$\frac{Q+({\bm{\beta}}+D_{f}-\hat{W})(p_{\max}+2\gamma)+{\bm{\beta}}\left(\bar{p}(c+\varepsilon)+2\delta\right)+(D_{f}-\hat{V}_{f})(p_{\max}(c+\varepsilon)+2\delta)-cp_{\min}\hat{W}}{\sum_{i=1}^{n}G_{i}(\beta_{i})+\varepsilon\bar{p}{\bm{\beta}}+\sum_{\tau=1}^{T}(1+\varepsilon)H_{\tau}(f_{\tau})+(D_{f}+{\bm{\beta}})\frac{2\kappa}{T}}>\alpha_{\texttt{T}}.$ (31)
@@
$=\frac{\sum_{i=1}^{n}\sum_{j=1}^{\nu_{i}}\left[X_{b}^{(i,j)}\right]+\sum_{\tau=1}^{T}\left[X_{f}^{(\tau)}\right]}{\sum_{i=1}^{n}\sum_{j=1}^{\nu_{i}}\left[\left(\hat{\phi}_{b}^{(i,j)}(\hat{w}_{b}^{(i,j)})\!-\!2\gamma\!+\!\varepsilon\bar{p}\!+\!\frac{2\kappa}{T}\right)B_{b}^{(i,j)}\right]\!+\!\sum_{\tau=1}^{T}\left[\frac{1+\varepsilon}{1+c+\varepsilon}\left(\hat{\phi}_{f}^{(\tau)}(\hat{w}_{f}^{(\tau)})\!+\!\hat{\psi}_{f}^{(\tau)}(\hat{v}_{f}^{(\tau)})\!-\!2\kappa\right)f_{\tau}\!+\!f_{\tau}\frac{2\kappa}{T}\right]},$ (32)
@@
G.2 PALD Robustness Certificate for OSDM-T
@@
We now proceed to prove the worst-case competitive ratio of PALD– for notational brevity, the following considers any arbitrary $\mathcal{I}\in\Omega$ (i.e., any arbitrary instance of OSDM). We begin the proof by stating a lower bound on the cost of the offline optimal solution OPT, leveraging the same logic as Lemma E.7.
@@
$\texttt{PALD}(\mathcal{I})$ $\leq\sum_{i=1}^{n}\sum_{j=1}^{\nu_{i}}\int_{0}^{\hat{w}_{b}^{(i,j)}}\kern-10.00002pt\hat{\phi}_{b}^{(i,j)}(u)du+\sum_{\tau=1}^{T}\left(\int_{0}^{\hat{w}_{f}^{(\tau)}}\kern-10.00002pt\hat{\phi}_{f}^{(\tau)}(u)du+\int_{0}^{\hat{v}_{f}^{(\tau)}}\kern-10.00002pt\hat{\psi}^{(\tau)}(z)dz\right)$ (33)
$\quad+\left(D_{b}+D_{f}-\sum_{i=1}^{n}\sum_{j=1}^{\nu_{i}}\hat{w}_{b}^{(i,j)}-\sum_{\tau=1}^{T}\hat{w}_{f}^{(\tau)}\right)(p_{\max}+2\eta)$ (34)
$\quad+D_{b}\left(\bar{p}(c+\varepsilon)+2\delta\right)+\left(D_{f}-\sum_{\tau=1}^{T}\hat{v}_{f}^{(\tau)}\right)(p_{\max}(c+\varepsilon)+2\delta)$ (35)
$\quad-cp_{\min}\left(\sum_{i=1}^{n}\sum_{j=1}^{\nu_{i}}\hat{w}_{b}^{(i,j)}+\sum_{\tau=1}^{T}\hat{w}_{f}^{(\tau)}\right)+\hat{s}p_{\max}.$ (36)
@@
The proof follows by the same logic as Lemma G.3 – the main change is with respect to (34), to capture the tracking cost (with parameter $\eta$ on the purchasing side of the problem). ∎
@@
$\frac{Q+({\bm{\beta}}+D_{f}-\hat{W})(p_{\max}+2\eta)+{\bm{\beta}}\left(\bar{p}(c+\varepsilon)+2\delta\right)+(D_{f}-\hat{V}_{f})(p_{\max}(c+\varepsilon)+2\delta)-cp_{\min}\hat{W}}{\sum_{i=1}^{n}G_{i}(\beta_{i})+\varepsilon\bar{p}{\bm{\beta}}+\sum_{\tau=1}^{T}(1+\varepsilon)H_{\tau}(f_{\tau})+(D_{f}+{\bm{\beta}})\frac{2\delta}{T}}>\alpha_{\texttt{T}}.$ (37)
@@
$=\frac{\sum_{i=1}^{n}\sum_{j=1}^{\nu_{i}}\left[X_{b}^{(i,j)}\right]+\sum_{\tau=1}^{T}\left[X_{f}^{(\tau)}\right]}{\sum_{i=1}^{n}\sum_{j=1}^{\nu_{i}}\left[\left(\hat{\phi}_{b}^{(i,j)}(\hat{w}_{b}^{(i,j)})\!-\!2\eta\!+\!\varepsilon\bar{p}\!+\!\frac{2\delta}{T}\right)B_{b}^{(i,j)}\right]\!+\!\sum_{\tau=1}^{T}\left[\frac{1+\varepsilon}{1+c+\varepsilon}\left(\hat{\phi}_{f}^{(\tau)}(\hat{w}_{f}^{(\tau)})\!+\!\hat{\psi}_{f}^{(\tau)}(\hat{v}_{f}^{(\tau)})\!-\!2(\eta+\delta)\right)f_{\tau}\!+\!f_{\tau}\frac{2\delta}{T}\right]},$ (38)

### v2

@@ Appendix F Deferred Proofs from § 5
To solve for $\alpha^{\star}$, we can use Grönwall’s inequality (Mitrinovic et al., 1991, Theorem 1, p. 356). First, observe that $h(x)$ must satisfy the following:
@@
By Grönwall’s inequality (Mitrinovic et al., 1991, Theorem 1, p. 356), it follows that:
@@
Recall that we have a boundary condition that $h(p_{\min})=1$—we can combine this with the above to obtain:
@@
$\alpha^{\star}\geq\frac{\omega}{\left[W\left(-\frac{\left((1+c+\varepsilon)p_{\max}-(1+\varepsilon)p_{\min}\right)e^{\frac{-\nicefrac{{\omega 2\kappa}}{{T}}-p_{\min}c-(1+c+\varepsilon)p_{\max}}{(1+c+\varepsilon)p_{\max}+2\kappa}}}{(1+c+\varepsilon)p_{\max}+2\kappa}\right)+\frac{(1+c+\varepsilon)p_{\max}+p_{\min}c-\nicefrac{{\omega 2\kappa}}{{T}}}{(1+c+\varepsilon)p_{\max}+2\kappa}\right]},$
@@
Let $h(x)$ denote a conversion function that maps $p_{\min},p_{\max}]\to[0,1]$—this describes ALG’s purchasing decisions before the final batch of prices ($p_{\max}$) in the instance. Since ALG is deterministic and purchasing is unidirectional (irrevocable), we must have that $h(x-\varrho)\geq h(x)$, i.e., $h(x)$ is non-increasing in $p_{\min},p_{\max}]$. Intuitively, the entire demand should be purchased before the end of the sequence if the lowest price appears, i.e., $h(p_{\min})=1$.
@@
By Grönwall’s inequality (Mitrinovic et al., 1991, Theorem 1, p. 356), it follows that:
@@
Recall that we have a boundary condition that $h(p_{\min})=1$—we can combine this with the above to obtain:
@@
(26) $\alpha_{\texttt{IF}}\geq\left[W\left(-\frac{\left(p_{\max}-p_{\min}\right)e^{-1}}{p_{\max}+\frac{2\kappa}{T(1+\varepsilon)}}\right)+1\right]^{-1}.$
@@
By Grönwall’s inequality (Mitrinovic et al., 1991, Theorem 1, p. 356), it follows that:
@@
Recall that we have a boundary condition that $h(p_{\min})=1$—we can combine this with the above to obtain:
@@
(27) $\alpha_{\texttt{IB}}\geq\left[W\left(-\frac{\left((1-c)p_{\max}-p_{\min}\right)e^{-\frac{(1+\varepsilon)p_{\max}+2\delta+\frac{2\gamma}{T}}{(1+\varepsilon)p_{\max}+2\kappa}}}{(1+\varepsilon)p_{\max}+2\kappa}\right)+\frac{(1+\varepsilon)p_{\max}+2\delta+\frac{2\gamma}{T}}{(1+\varepsilon)p_{\max}+2\kappa}\right]^{-1}.$
@@
Let $h(x)$ denote a conversion function that maps $p_{\min},p_{\max}]\to[0,1]$—this describes ALG’s purchasing decisions before the final batch of prices ($p_{\max}$) in the instance. Since ALG is deterministic and purchasing is unidirectional (irrevocable), we must have that $h(x-\varrho)\geq h(x)$, i.e., $h(x)$ is non-increasing in $p_{\min},p_{\max}]$. Intuitively, the entire demand should be purchased before the end of the sequence if the lowest price appears, i.e., $h(p_{\min})=1$.
@@
By Grönwall’s inequality (Mitrinovic et al., 1991, Theorem 1, p. 356), it follows that:
@@
Recall that we have a boundary condition that $h(p_{\min})=1$—we can combine this with the above to obtain:
@@
Due to the tracking target, any procurement before the second-to-last batch incurs a tracking penalty of $\eta h(x)$, since the target is zero during this time. Furthermore, during the last two batches, ALG incurs a tracking penalty of $\eta+\eta(1-h(x))$, since the target is $\nicefrac{{1}}{{m}}$ during the first $m$ prices in the batch, and zero during the second $m$ prices in the batch—note that ALG must purchase (1h(x))(1-h(x)) at the last price to satisfy the deadline. Thus, the total tracking penalty incurred by ALG is given by $2\eta$.
@@
To solve for $\alpha^{\star}$, we can use Grönwall’s inequality (Mitrinovic et al., 1991, Theorem 1, p. 356). First, observe that $h(x)$ must satisfy the following:
@@
Recall that we have a boundary condition that $h(p_{\min})=1$—we can combine this with the above to obtain:
@@
$\alpha^{\star}\geq\frac{\omega}{\left[W\left(-\frac{\left((1+c+\varepsilon)p_{\max}-(1+\varepsilon)p_{\min}\right)e^{\frac{-\nicefrac{{\omega 2\delta}}{{T}}-p_{\min}c-(1+c+\varepsilon)p_{\max}}{(1+c+\varepsilon)p_{\max}+2(\eta+\delta)}}}{(1+c+\varepsilon)p_{\max}+2(\eta+\delta)}\right)+\frac{(1+c+\varepsilon)p_{\max}+p_{\min}c-\nicefrac{{\omega 2\delta}}{{T}}}{(1+c+\varepsilon)p_{\max}+2(\eta+\delta)}\right]},$
@@
We now proceed to prove the worst-case competitive ratio of PALD—for notational brevity, the following considers any arbitrary $\mathcal{I}\in\Omega$ (i.e., any arbitrary instance of OSDM). We recall the definition of active and inactive periods from Definition E.2, and the additional notations relating to driver indexing used in the proof of Theorem 4.3 in Appendix E.
@@
(28) $\texttt{PALD}(\mathcal{I})$ $\leq\sum_{i=1}^{n}\sum_{j=1}^{\nu_{i}}\int_{0}^{\hat{w}_{b}^{(i,j)}}\kern-10.00002pt\hat{\phi}_{b}^{(i,j)}(u)du+\sum_{\tau=1}^{T}\left(\int_{0}^{\hat{w}_{f}^{(\tau)}}\kern-10.00002pt\hat{\phi}_{f}^{(\tau)}(u)du+\int_{0}^{\hat{v}_{f}^{(\tau)}}\kern-10.00002pt\hat{\psi}^{(\tau)}(z)dz\right)$
(29) $\quad+\left(D_{b}+D_{f}-\sum_{i=1}^{n}\sum_{j=1}^{\nu_{i}}\hat{w}_{b}^{(i,j)}-\sum_{\tau=1}^{T}\hat{w}_{f}^{(\tau)}\right)(p_{\max}+2\gamma)$
(30) $\quad+D_{b}\left(\bar{p}(c+\varepsilon)+2\delta\right)+\left(D_{f}-\sum_{\tau=1}^{T}\hat{v}_{f}^{(\tau)}\right)(p_{\max}(c+\varepsilon)+2\delta)$
(31) $\quad-cp_{\min}\left(\sum_{i=1}^{n}\sum_{j=1}^{\nu_{i}}\hat{w}_{b}^{(i,j)}+\sum_{\tau=1}^{T}\hat{w}_{f}^{(\tau)}\right)+\hat{s}p_{\max}.$
@@
The proof follows by the same logic as Lemma E.4—the main change is with respect to (28), to capture the fact that the threshold functions used by PALD are the learned variants, denoted by $\hat{\phi}_{b}$ and $(\hat{\phi}_{f},\hat{\psi}_{f})$, instead of the analytical variants $\phi_{b}$ and $(\phi_{f},\psi_{f})$ used by PAAD.
@@
(32) $\frac{Q+({\bm{\beta}}+D_{f}-\hat{W})(p_{\max}+2\gamma)+{\bm{\beta}}\left(\bar{p}(c+\varepsilon)+2\delta\right)+(D_{f}-\hat{V}_{f})(p_{\max}(c+\varepsilon)+2\delta)-cp_{\min}\hat{W}}{\sum_{i=1}^{n}G_{i}(\beta_{i})+\varepsilon\bar{p}{\bm{\beta}}+\sum_{\tau=1}^{T}(1+\varepsilon)H_{\tau}(f_{\tau})+(D_{f}+{\bm{\beta}})\frac{2\kappa}{T}}>\alpha_{\texttt{T}}.$
@@
(33) $=\frac{\sum_{i=1}^{n}\sum_{j=1}^{\nu_{i}}\left[X_{b}^{(i,j)}\right]+\sum_{\tau=1}^{T}\left[X_{f}^{(\tau)}\right]}{\sum_{i=1}^{n}\sum_{j=1}^{\nu_{i}}\left[\left(\hat{\phi}_{b}^{(i,j)}(\hat{w}_{b}^{(i,j)})\!-\!2\gamma\!+\!\varepsilon\bar{p}\!+\!\frac{2\kappa}{T}\right)B_{b}^{(i,j)}\right]\!+\!\sum_{\tau=1}^{T}\left[\frac{1+\varepsilon}{1+c+\varepsilon}\left(\hat{\phi}_{f}^{(\tau)}(\hat{w}_{f}^{(\tau)})\!+\!\hat{\psi}_{f}^{(\tau)}(\hat{v}_{f}^{(\tau)})\!-\!2\kappa\right)f_{\tau}\!+\!f_{\tau}\frac{2\kappa}{T}\right]},$
@@
G.2. PALD Robustness Certificate for OSDM-T
@@
We now proceed to prove the worst-case competitive ratio of PALD—for notational brevity, the following considers any arbitrary $\mathcal{I}\in\Omega$ (i.e., any arbitrary instance of OSDM). We begin the proof by stating a lower bound on the cost of the offline optimal solution OPT, leveraging the same logic as Lemma E.7.
@@
(34) $\texttt{PALD}(\mathcal{I})$ $\leq\sum_{i=1}^{n}\sum_{j=1}^{\nu_{i}}\int_{0}^{\hat{w}_{b}^{(i,j)}}\kern-10.00002pt\hat{\phi}_{b}^{(i,j)}(u)du+\sum_{\tau=1}^{T}\left(\int_{0}^{\hat{w}_{f}^{(\tau)}}\kern-10.00002pt\hat{\phi}_{f}^{(\tau)}(u)du+\int_{0}^{\hat{v}_{f}^{(\tau)}}\kern-10.00002pt\hat{\psi}^{(\tau)}(z)dz\right)$
(35) $\quad+\left(D_{b}+D_{f}-\sum_{i=1}^{n}\sum_{j=1}^{\nu_{i}}\hat{w}_{b}^{(i,j)}-\sum_{\tau=1}^{T}\hat{w}_{f}^{(\tau)}\right)(p_{\max}+2\eta)$
(36) $\quad+D_{b}\left(\bar{p}(c+\varepsilon)+2\delta\right)+\left(D_{f}-\sum_{\tau=1}^{T}\hat{v}_{f}^{(\tau)}\right)(p_{\max}(c+\varepsilon)+2\delta)$
(37) $\quad-cp_{\min}\left(\sum_{i=1}^{n}\sum_{j=1}^{\nu_{i}}\hat{w}_{b}^{(i,j)}+\sum_{\tau=1}^{T}\hat{w}_{f}^{(\tau)}\right)+\hat{s}p_{\max}.$
@@
The proof follows by the same logic as Lemma G.3—the main change is with respect to (35), to capture the tracking cost (with parameter $\eta$ on the purchasing side of the problem). ∎
@@
(38) $\frac{Q+({\bm{\beta}}+D_{f}-\hat{W})(p_{\max}+2\eta)+{\bm{\beta}}\left(\bar{p}(c+\varepsilon)+2\delta\right)+(D_{f}-\hat{V}_{f})(p_{\max}(c+\varepsilon)+2\delta)-cp_{\min}\hat{W}}{\sum_{i=1}^{n}G_{i}(\beta_{i})+\varepsilon\bar{p}{\bm{\beta}}+\sum_{\tau=1}^{T}(1+\varepsilon)H_{\tau}(f_{\tau})+(D_{f}+{\bm{\beta}})\frac{2\delta}{T}}>\alpha_{\texttt{T}}.$
@@
(39) $=\frac{\sum_{i=1}^{n}\sum_{j=1}^{\nu_{i}}\left[X_{b}^{(i,j)}\right]+\sum_{\tau=1}^{T}\left[X_{f}^{(\tau)}\right]}{\sum_{i=1}^{n}\sum_{j=1}^{\nu_{i}}\left[\left(\hat{\phi}_{b}^{(i,j)}(\hat{w}_{b}^{(i,j)})\!-\!2\eta\!+\!\varepsilon\bar{p}\!+\!\frac{2\delta}{T}\right)B_{b}^{(i,j)}\right]\!+\!\sum_{\tau=1}^{T}\left[\frac{1+\varepsilon}{1+c+\varepsilon}\left(\hat{\phi}_{f}^{(\tau)}(\hat{w}_{f}^{(\tau)})\!+\!\hat{\psi}_{f}^{(\tau)}(\hat{v}_{f}^{(\tau)})\!-\!2(\eta+\delta)\right)f_{\tau}\!+\!f_{\tau}\frac{2\delta}{T}\right]},$